OpenStudy (anonymous):

A typical car engine produces 44,000 kJ of heat in an hour. The cooling system has a capacity of 8.40 L and is filled with a 50:50 mixture of anti-freeze (specific heat capacity of 8.37 J/g-oC, density of 1.038 g/mL). In maintaining a constant engine temperature, the cooling system temperature increased from 95oC to 110oC. (Assume this is within allowable limits.) How many times did the cooling system turn over its volume of anti-freeze to maintain the engine temperature during its hour of operation? Show your work.

OpenStudy (anonymous):

My work so far: Q = mct 44000 = m*8.37*15 M = 350.457985 D=m/v 1.038 = 350.457985/v Mass/d = volume 350.457985/1.038 = 337.62811657ml 0.3376281L

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I don't know how to use this information to figure out the answer.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@wio @dan815

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@freckles